Light of wavelength λ = 600 nm shines through two narrow slits which are 660 μm apart. What is the maximum number of interference maxima which could conceivably be observed (assuming that diffraction minima do not extinguish them and the screen is arbitrarily large)?
Answer #1

Solution :

dsinθ=nλ

dsin⁡θ/λ = n

For maximum value of interference maxima,

sinθ = 1

n = 660 × 10-6 ×1 / (600 × 10-9)

n = 1100.

So, total number of interference maxima = 2n + 1 = 2201.